OpenStudy (anonymous):

Find the equation of the line and write in general form. a) y-intercept = -2 and x - intercept = 3 b) Through (2, 5) and (3, 6) c) Line parallel to the line 4x + 2y = 8 through (0, 4)

OpenStudy (bradely):

a) x/a+y/b =1 x/3+y/-2=1 2x-3y=6 Source: http://www.mathskey.com/question2answer/

OpenStudy (bradely):

b) Slope m =(6-5)/(3-2) =1 y=x+c 5=2+c c=3 y=x+3 x-y+3=0 Source: http://www.mathskey.com/question2answer/